Cryptology Problems

To be discussed Wednesday, February 18, 1998.

1. This was encrypted using a Caesar cipher. Decipher it. YMNX HWDUYTLWFR NX JFXD YT IJHNUMJW.

2. This was entitled ”Cold Country”. It was encrypted using a monoalphabetic substiution cipher. Decipher it. TOWWJPHJC ZY RXW PHOTWYR ZYPHJC ZJ RXW SFOPC. UFYR FB ZR ZY QFIWOWC SZRX ZQW RXFMYHJCY FB BWWR CWWD.

3.Suppose you had two examples of ciphertext, both enciphered using periodic polyalphabetic ciphers. How would you make an intelligent guess as to whether or not the same sequence of substitution alphabets was used, without making any attempt at deciphering? Is the assumption that the ciphers are periodic necessary?

4. Give a linear feedback shift register over the field GF(2) which will produce the sequence which is 0 1 1 0 repeated over and over again, indefinitely. (Give the shortest one you can find.)

5. Suppose that you are able to obtain the cyphertext “01101010101101”, and you learn that the first eight bits of the plaintext are “10111011”. You know that the encryption was done with the aid of a linear feedback shift register over the field GF(2), with length four. Determine the linear feedback shift register and the remainder of the message.

6. Use the output of the following linear feedback shift register as a onetime pad to encode the message “0110101000110”. The register of the linear feedback shift register initially contains (0,1,1,1), and the tap sequence is (0,1,0,1).

a What is the cryptotext?

b What is the length of the period of the sequence generated by the linear feedback shift register?